OpenStudy (anonymous):

1/n{(1/n)^9+(2/n)^9+(3/n)^9+⋯+(n/n)^9}

OpenStudy (loser66):

Can someone type it in latex, please. I don't get the problem.

OpenStudy (anonymous):

Is this your problem? \[\frac{ 1 }{ n }\left\{ \left( \frac{ 1 }{ n } \right)^9+\left( \frac{ 2 }{ n } \right)^9+\left( \frac{ 3 }{ n } \right)^9+...+\left( \frac{ n }{ n } \right)^9 \right\}\]

OpenStudy (mathmath333):

^

OpenStudy (mathmath333):

but the asker didnt write what to do with it

OpenStudy (anonymous):

yes ,it is not clear ,what he wants to do.

OpenStudy (zarkon):

probably wants the limit as \(n\to\infty\)